A new brick wall has been formed in between the ground floor hall and the staircase/lift space. The wall goes all the way to the top of the building. This was the access to the old cupboard in the downstairs ‘interview room’.
The heating pipes in the ground floor ceiling have been lagged.
New stones have been installed in the arch beneath the old first-floor fire escape.
Pad stones being installed, starting today for the new staircase support.
The ceiling repairs have been finished to the curved sections of first-floor ceiling.
The internal cornice has been repaired where the old lights were.
Paint stripping to wooden panels is ongoing.
The plasterers have now started to repair walls in the proposed kitchen area. This was the old office where the large marble tablets listing the mayors used to hang.
The south and west side have been slated completely. The East & North sides are ongoing.
This is the east elevation.
New leadwork being formed for the clock-tower.
New timber mouldings have been installed on the cupola.
New lead skirts to the bottom of each cupola column have been installed.
New heating pipes installed in the attic floor void.
Proposed bath waste pipe started.
All old wiring and electrical equipment are being removed.
Also while stripping the attic, some names were found dating from previous restorations…. does anyone know Nobby or D Baly?